Help I'm lost. PLEASE!
An annuity has a record of 9.58% annual growth compounded annually. How long will it take to double your money in this fund

Answer by stanbon(75887)

An annuity has a record of 9.58% annual growth compounded annually. How long will it take to double your money in this fund
-------------
Formula:
A=P(1+r/n)^(nt)
A is the future value
P is the money you start with
r is the annual rate of interest
n is the number of times you compound in a year
t is the number of years you compound
---------
In your problem you want the future value to be twice the start value:
i.e. you want A=2P
2P=P(1+0.0958/1)^(1*t)
2=(1.0958)^t
Take the log of both sides to get:
log2 = t (log 1.0958)
0.30102999... = t(0.0397329...)
t=7.58 years
The money will double in 7.58 years
Cheers,
Stan H.